# Smooth Sailing: Lipschitz-Driven Uncertainty Quantification for Spatial Association

### v2

@@ 3 Lipschitz-Driven Inference
For $\mathcal{S}$ as in Proposition 8, with the assumptions and notation of Theorem 7, but with $\sigma^{2}$ unknown, the confidence interval $I^{\Psi}$ from Algorithm 1 has asymptotic (with fixed $M$ and as $N\to\infty$) coverage at level $1-\alpha$.
Choice of the Lipschitz Constant $L$. The Lipschitz assumption allows us to make inferences about the target data from the source data. The Lipschitz constant subsequently enters our intervals in two principal ways: via $B$ and via $\hat{\sigma}^{2}_{N}$. Intuitively, larger values of $L$ (allowing for less smooth responses) lead to our algorithm constructing confidence intervals with larger bounds ($B$) on the bias but smaller estimated residual variance ($\hat{\sigma}^{2}_{N}$). We depict this trade-off in a concrete example in Section 4 (Fig. 2).
Ultimately the choice of Lipschitz constant must be guided by domain knowledge. We give a concrete example describing our choice of the Lipschitz constant in our real-data experiment on tree cover (Section 4). In our simulated experiments, we know the minimum value for which Assumption 4 holds; call it $L_{0}$. So we first choose $L=L_{0}$. Then we perform ablation studies in both simulated and real data showing that we essentially maintain coverage while varying $L$ over an order of magnitude around our initial choices. We show that further decreasing $L$ can decrease coverage and discuss why it is useful to err on the side of conservatism (i.e., a larger $L$). However, we expect even small values of $L$ to improve upon classical confidence intervals in terms of coverage, since classical confidence intervals do not account for bias at all; our method similarly ignores bias when $L=0$.
@@ 4 Experiments
Figure 1: Coverages (left) and confidence interval widths (right) for our method as well as 5 other methods (3 methods in the lower experiment). In the upper experiment, our method and GP BCIs consistently achieve the nominal coverage (95%); the GP BCIs line (dashed blue) overlaps with ours (solid black) for most shifts. Of the two methods with correct coverage, our method yields much narrower intervals. In the lower experiment, only our method achieves the nominal coverage. The shaded region for coverage is a (conservative) 95% confidence interval while the shaded region for CI width is ±$\pm 2$ standard deviations; for more detail, see Section C.1.
@@
Ordinary Least Squares (OLS). We treat the noise variance as unknown and estimate it as the average squared residual, with a correction for the number of degrees of freedom and a $t$-statistic instead of a $z$-statistic [22, pp. 50–52].
Sandwich Estimator. The sandwich estimator [27, 71, 72] uses the same point estimate as OLS but a different variance estimate. We take the variance estimate from MacKinnon and White, [36, Equation 6]: $\frac{1}{N-P}e_{p}^{\mathrm{T}}(X^{\mathrm{T}}X)^{-1}(X^{\mathrm{T}}RX)(X^{\mathrm{T}}X)^{-1}e_{p}$, where $R$ is a diagonal matrix with the squared residuals as entries.
Importance Weighted Least Squares (KDEIW). As suggested by Shimodaira, [57, Section 9], we calculate importance weights via kernel density estimation (KDE). We select the bandwidth parameter with $5$-fold cross validation; see Section B.7. Given the KDE weights, we use the point estimate and confidence interval from weighted least squares.
@@
Gaussian Process Bayesian Credible Intervals (GP BCIs). We use the model $Y(S)=\theta^{\mathrm{T}}X(S)+g(S)+\epsilon,$ with $\theta^{\mathrm{T}}\sim\mathcal{N}(0,\lambda^{2}I_{P})$, $g\sim\mathcal{GP}(0,k_{\gamma})$, $k_{\gamma}$ an isotropic Matérn 3/2 kernel function with hyperparameters $\gamma$, and $\epsilon\sim\mathcal{N}(0,\delta^{2})$. We select $\{\lambda,\gamma,\delta\}$ by maximum likelihood. We report posterior credible intervals for $\theta_{p}$.
Single Covariate Simulation. In our first simulation, the source locations are uniform on $\mathcal{S}=[-1,1]^{2}$ (blue points in Fig. 4 left plots). The target locations are uniform on $[\frac{-1+\mathrm{shift}}{1+|\mathrm{shift}|},\frac{1+\mathrm{shift}}{1+|\mathrm{shift}|}]$, where $\mathrm{shift}$ controls the degree of distribution shift between source and target (orange points in Fig. 4 left plots). In this experiment, the single covariate $X=\chi(S)=S^{(1)}+S^{(2)}$ (Fig. 4, third plot). And the response is $Y=X+\frac{1}{2}((S^{(1)})^{2}+(S^{(2)})^{2})+\epsilon$, with $\epsilon\sim\mathcal{N}(0,0.1^{2})$. Fig. 4, fourth plot, shows the conditional expectation of the response given location. We can compute the ground truth parameter in closed form because we have access to the conditional expectation of the response (Eq. 4). We vary $\mathrm{shift}\in[0,\pm 0.2,\pm 0.4,\pm 0.6,\pm 0.8]$ and run 250 seeds for each $\mathrm{shift}$.
Figure 1, top left, shows that only our method and the GP consistently achieve nominal coverage. Given correct coverage, narrower (i.e., more precise) confidence intervals are desirable; Fig. 1, top right, shows that our method yields narrower intervals than the GP. KDEIW comes close to achieving nominal coverage when there is no shift. But under shift with limited data, it is not able to fully debias the estimate, and coverage drops. For large $M$ (here $M=100$), we expect the sandwich estimator to achieve nominal coverage at $\mathrm{shift}=0$ since it is guaranteed to cover the population under first-order misspecification without distribution shift [27, 72]. But under any of the depicted non-zero shifts, the sandwich, OLS, and GLS achieve zero coverage.
@@
The extreme narrowness of the OLS, sandwich, KDEIW, and GLS intervals (Fig. 1, top right) suggests that the problem with these methods is exactly their overconfidence. Because these approaches assume that the estimator is unbiased (or can be debiased), and that errors are independent and Gaussian, their intervals contract far too quickly, even with small amounts of data (here, $N=300$). Essentially, these methods’ reliance on strong modeling assumptions leads to non-robust coverage.
We note that the confidence intervals and coverages for each method have (approximately) the same values for either ±$\pm\mathrm{shift}$ in this experiment (Fig. 1, top right) because the covariate is symmetric around the line $S^{(1)}=S^{(2)}$; see the middle right plot of Fig. 4 in Appendix C. So positive and negative $\mathrm{shift}$ values move the source and target locations symmetrically across the $S^{(1)}=S^{(2)}$ line. We do not expect such a symmetry for general covariates and will not see it in our next simulation (Fig. 1, bottom right).
Simulation with Several Covariates. Our second simulation generates locations as in the previous experiment. Now we use $N=10{,}000$ and $M=100$. And we generate 3 covariates, $X^{(1)}=\sin(S^{(1)})+\cos(S^{(2)}),$ $X^{(2)}=\cos(S^{(1)})-\sin(S^{(2)})$ and $X^{(3)}=S^{(1)}+S_{2}$. The response is $Y=X^{(1)}X^{(2)}+\frac{1}{2}((S^{(1)})^{2}+(S^{(2)})^{2})+\epsilon$ with $\epsilon\sim\mathcal{N}(0,0.1^{2})$. We focus on inference for the first coefficient. Calculation of $\hat{\sigma}_{N}^{2}$ scales poorly with $N$ due to needing to solve a quadratic program. So here we instead estimate $\sigma^{2}$ using the squared error of leave-one-out $1$-nearest neighbor regression fit on the source data. See Section B.6 for details. We compare against the same set of methods except we do not include GLS or GP BCIs since these require further approximations to scale for this $N$.
We again find that our method achieves coverage while the other methods do not (Fig. 1, bottom left). In this experiment, no other method achieves coverage over 30% across any $\mathrm{shift}$ value (even 0). As before, competing methods are overconfident, with very small CI widths (Fig. 1, bottom right).
For methods besides our own, coverage levels at 00 are generally lower in this experiment at $\mathrm{shift}=0$ than in the previous experiment. The difference is that $N$ is much larger here (making confidence intervals narrower and exacerbating overconfidence), while $M$ is the same. The sandwich estimator covers the analogue of $\theta^{\star}_{\text{OLS},p}$ where spatial locations are treated as random. So it has good coverage at $\mathrm{shift}=0$ when $M\gg N$, but not when $M\ll N$.
@@
Effect of Lipschitz Constant on Confidence Intervals in Simulation Experiment. Above, we know and use the minimum value ($L_{0}$) for which Assumption 4 holds; that is, $L=L_{0}$ ($L_{0}=2\sqrt{2}$ and $3\sqrt{2}$, respectively). Now we repeat the first simulation but vary $L\in\{0.1,0.5,1.0,2.0,3.5,5,7.5,10\}$. All of these $L$ values yield coverage of $1.0$ for our method, above the nominal coverage of $0.95$, even though coverage is not guaranteed by our theory for $L\leq 2\sqrt{2}\approx 2.8$.
@@
We next show that the confidence interval width reflects a bias-randomness trade-off as $L$ varies. If the noise were known, the confidence interval would be monotonically increasing in $L$. Since the noise is unknown, only the bias contribution to the interval width ($2B$, Step 4 in Algorithm 1) increases (Fig. 2, middle and left, $\times$). Conversely, smaller values of $L$ yield larger values for $\hat{\sigma}^{2}_{N}$, so the randomness contribution to the interval width ($2c$, Step 5 in Algorithm 1) increases (Fig. 2, middle and left, $\square$). The full confidence interval width, $2B+2c\Delta(\alpha)$, is not monotonic in $L$.
[…]
@@
We compare coverage for confidence intervals of the three parameters, $\theta_{\text{Aridity\_Index}},\theta_{\text{Elevation}},\text{ and }\theta_{\text{Slope}}$. We discuss how we evaluate coverage in Section D.3. In the top row of Fig. 3, we see that our method is the only one to achieve the $95\%$ nominal coverage for all three parameters. Conversely, for the Slope parameter, every other method achieves coverage at most $54\%$. In the bottom row of Fig. 10, we again see that alternative methods fail to provide coverage due to their overconfidence (small widths); see also Fig. 10, which shows all methods’ constructed confidence intervals across all three parameters for 7 of the 250 seeds. In Section D.5, we conduct a similar analysis but with target locations in the Southeast, rather than West, of CONUS. The results align with our discussion here.
@@
Choice of Lipschitz Constant in the Tree Cover Experiment. For the tree cover experiment, we leverage domain knowledge to set the Lipschitz constant to $L=0.2$, in units of percent tree cover per kilometer (km). This choice implies that a $1\%$ change in tree cover corresponds to moving $1/0.2=5$ km. To arrive at this choice, we observe that in certain regions of the U.S., such as the Midwest, tree coverage remains relatively uniform over several kilometers, so smaller Lipschitz constants (e.g., $L=0.02$, corresponding to a $1\%$ change over 50 km) would be appropriate. However, in other regions — such as the western U.S., where elevation changes are more pronounced (e.g., the Rockies, California, and the Pacific Northwest) — tree cover can change sharply over short distances. To account for these variations conservatively, we choose $L=0.2$. More generally, for real-world applications, we recommend the following strategy: (i) use domain knowledge to select a reasonable Lipschitz constant for the response variable, and (ii) inflate the Lipschitz constant to ensure a conservative estimate (which is more likely to satisfy Assumption 4).
@@
Effect of Lipschitz Constant on Confidence Intervals in Tree Cover Experiment. In Fig. 9, we show the coverage and width of our confidence intervals across three orders of magnitude of Lipschitz constants ($L$ from 0.001 to 1). For $L$ varying between 0.1 and 1 (a single order of magnitude variation around our chosen value of 0.2), we find that coverage is always met except in one case (Aridity Index with $L=0.1$), where it is very close to nominal (89% instead of 95%). For Slope and Elevation, coverage is met or very nearly met for all $L$ values. For Aridity Index, coverage is low for $L\leq 0.1$. Meanwhile, confidence intervals become noticeably wider for $L>0.5$ while remaining relatively stable for smaller values. These results support our intuition to err on the side of larger $L$ values to be conservative and maintain coverage.

### v3

@@ 3 Lipschitz-Driven Inference
For $\mathcal{S}$ as in Proposition˜8, with the assumptions and notation of Theorem˜7, but with $\sigma^{2}$ unknown, the confidence interval $I^{\Psi}$ from Algorithm˜1 has asymptotic (with fixed $M$ and as $N\to\infty$) coverage at level $1-\alpha$.
Choice of the Lipschitz Constant $L$. The Lipschitz assumption allows us to make inferences about the target data from the source data. The Lipschitz constant subsequently enters our intervals in two principal ways: via $B$ and via $\hat{\sigma}^{2}_{N}$. Intuitively, larger values of $L$ (allowing for less smooth responses) lead to our algorithm constructing confidence intervals with larger bounds ($B$) on the bias but smaller estimated residual variance ($\hat{\sigma}^{2}_{N}$). We depict this trade-off in a concrete example in Section˜4 (Fig.˜2).
Ultimately the choice of Lipschitz constant must be guided by domain knowledge. We give one concrete example describing our choice of the Lipschitz constant in our real-data experiment on tree cover (Section˜4). We give a second concrete example of how to select the Lipschitz constant in Section˜B.3; in this case, the response is annual average PM2.5 over California. In our simulated experiments, we know the minimum value for which Assumption˜4 holds; call it $L_{0}$. So we first choose $L=L_{0}$. Then we perform ablation studies in both simulated and real data showing that we essentially maintain coverage while varying $L$ over an order of magnitude around our initial choices. We show that further decreasing $L$ can decrease coverage and discuss why it is useful to err on the side of conservatism (i.e., a larger $L$). However, we expect even small values of $L$ to improve upon classical confidence intervals in terms of coverage, since classical confidence intervals do not account for bias at all; our method similarly ignores bias when $L=0$.
@@ 4 Experiments
Figure 1: Coverages (left) and confidence interval widths (right) for our method as well as 5 other methods (3 methods in the lower experiment). In the upper experiment, our method and GP BCIs consistently achieve the nominal coverage (95%); the GP BCIs line (dashed blue) overlaps with ours (solid black) for most shifts. Of the two methods with correct coverage, our method yields much narrower intervals. In the lower experiment, only our method achieves the nominal coverage. The shaded region for coverage is a (conservative) 95% confidence interval while the shaded region for CI width is $\pm 2$ standard deviations; for more detail, see Section˜D.1.
@@
Ordinary Least Squares (OLS). We treat the noise variance as unknown and estimate it as the average squared residual, with a correction for the number of degrees of freedom and a $t$-statistic instead of a $z$-statistic [24, pp. 50–52].
Sandwich Estimator. The sandwich estimator [30, 77, 78] uses the same point estimate as OLS but a different variance estimate. We take the variance estimate from MacKinnon and White, [39, Equation 6]: $\frac{1}{N-P}e_{p}^{\mathrm{T}}(X^{\mathrm{T}}X)^{-1}(X^{\mathrm{T}}RX)(X^{\mathrm{T}}X)^{-1}e_{p}$, where $R$ is a diagonal matrix with the squared residuals as entries.
Importance Weighted Least Squares (KDEIW). As suggested by Shimodaira, [62, Section 9], we calculate importance weights via kernel density estimation (KDE). We select the bandwidth parameter with $5$-fold cross validation; see Section˜B.9. Given the KDE weights, we use the point estimate and confidence interval from weighted least squares.
@@
Gaussian Process Bayesian Credible Intervals (GP BCIs). We use the model $Y(S)=\theta^{\mathrm{T}}X(S)+h(S)+\epsilon,$ with $\theta^{\mathrm{T}}\sim\mathcal{N}(0,\lambda^{2}I_{P})$, $h\sim\mathcal{GP}(0,k_{\gamma})$, $k_{\gamma}$ an isotropic Matérn 3/2 kernel function with hyperparameters $\gamma$, and $\epsilon\sim\mathcal{N}(0,\delta^{2})$. We select $\{\lambda,\gamma,\delta\}$ by maximum likelihood. We report posterior credible intervals for $\theta_{p}$.
Single Covariate Simulation. In our first simulation, the source locations are uniform on $\mathcal{S}=[-1,1]^{2}$ (blue points in Fig.˜4 left plots). The target locations are uniform on $\frac{-1+\mathrm{shift}}{1+|\mathrm{shift}|},\frac{1+\mathrm{shift}}{1+|\mathrm{shift}|}]$, where $\mathrm{shift}$ controls the degree of distribution shift between source and target (orange points in Fig.˜4 left plots). In this experiment, the single covariate $X=\chi(S)=S^{(1)}+S^{(2)}$ (Fig.˜4, third plot). And the response is $Y=X+\frac{1}{2}((S^{(1)})^{2}+(S^{(2)})^{2})+\epsilon$, with $\epsilon\sim\mathcal{N}(0,0.1^{2})$. Fig.˜4, fourth plot, shows the conditional expectation of the response given location. We can compute the ground truth parameter in closed form because we have access to the conditional expectation of the response (Eq.˜4). We vary $\mathrm{shift}\in[0,\pm 0.2,\pm 0.4,\pm 0.6,\pm 0.8]$ and run 250 seeds for each $\mathrm{shift}$.
Figure˜1, top left, shows that only our method and the GP consistently achieve nominal coverage. Given correct coverage, narrower (i.e., more precise) confidence intervals are desirable; Fig.˜1, top right, shows that our method yields narrower intervals than the GP. KDEIW comes close to achieving nominal coverage when there is no shift. But under shift with limited data, it is not able to fully debias the estimate, and coverage drops. For large $M$ (here $M=100$), we expect the sandwich estimator to achieve nominal coverage at $\mathrm{shift}=0$ since it is guaranteed to cover the population under first-order misspecification without distribution shift [30, 78]. But under any of the depicted non-zero shifts, the sandwich, OLS, and GLS achieve zero coverage.
@@
The extreme narrowness of the OLS, sandwich, KDEIW, and GLS intervals (Fig.˜1, top right) suggests that the problem with these methods is exactly their overconfidence. Because these approaches assume that the estimator is unbiased (or can be debiased), and that errors are independent and Gaussian, their intervals contract far too quickly, even with small amounts of data (here, $N=300$). Essentially, these methods’ reliance on strong modeling assumptions leads to non-robust coverage.
We note that the confidence intervals and coverages for each method have (approximately) the same values for either $\pm\mathrm{shift}$ in this experiment (Fig.˜1, top right) because the covariate is symmetric around the line $S^{(1)}=S^{(2)}$; see the middle right plot of Fig.˜4 in Appendix˜D. So positive and negative $\mathrm{shift}$ values move the source and target locations symmetrically across the $S^{(1)}=S^{(2)}$ line. We do not expect such a symmetry for general covariates and will not see it in our next simulation (Fig.˜1, bottom right).
Simulation with Several Covariates. Our second simulation generates locations as in the previous experiment. Now we use N=10, [^0] N=10{,}000 and $M=100$. And we generate 3 covariates, $X^{(1)}=\sin(S^{(1)})+\cos(S^{(2)}),$ $X^{(2)}=\cos(S^{(1)})-\sin(S^{(2)})$ and $X^{(3)}=S^{(1)}+S_{2}$. The response is $Y=X^{(1)}X^{(2)}+\frac{1}{2}((S^{(1)})^{2}+(S^{(2)})^{2})+\epsilon$ with $\epsilon\sim\mathcal{N}(0,0.1^{2})$. We focus on inference for the first coefficient. Calculation of $\hat{\sigma}_{N}^{2}$ scales poorly with $N$ due to needing to solve a quadratic program. So here we instead estimate $\sigma^{2}$ using the squared error of leave-one-out $1$-nearest neighbor regression fit on the source data. See Section˜B.8 for details. We compare against the same set of methods except we do not include GLS or GP BCIs since these require further approximations to scale for this $N$.
We again find that our method achieves coverage while the other methods do not (Fig.˜1, bottom left). In this experiment, no other method achieves coverage over 30% across any $\mathrm{shift}$ value (even 0). As before, competing methods are overconfident, with very small CI widths (Fig.˜1, bottom right).
For methods besides our own, coverage levels at 0 are generally lower in this experiment at $\mathrm{shift}=0$ than in the previous experiment. The difference is that $N$ is much larger here (making confidence intervals narrower and exacerbating overconfidence), while $M$ is the same. The sandwich estimator covers the analogue of $\theta^{\star}_{\text{OLS},p}$ where spatial locations are treated as random. So it has good coverage at $\mathrm{shift}=0$ when $M\gg N$, but not when $M\ll N$.
@@
Effect of Lipschitz Constant on Confidence Intervals in Simulation Experiment. Above, we know and use the minimum value ($L_{0}$) for which Assumption˜4 holds; that is, $L=L_{0}$ ($L_{0}=2\sqrt{2}$ and $3\sqrt{2}$, respectively). Now we repeat the first simulation but vary $L\in\{0.1,0.5,1.0,2.0,3.5,5,7.5,10\}$. All of these $L$ values yield coverage of $1.0$ for our method, above the nominal coverage of $0.95$, even though coverage is not guaranteed by our theory for $L\leq 2\sqrt{2}\approx 2.8$.
@@
We next show that the confidence interval width reflects a bias-randomness trade-off as $L$ varies. If the noise were known, the confidence interval would be monotonically increasing in $L$. Since the noise is unknown, only the bias contribution to the interval width ($2B$, Step 4 in Algorithm˜1) increases (Fig.˜2, middle and left, $\times$). Conversely, smaller values of $L$ yield larger values for $\hat{\sigma}^{2}_{N}$, so the randomness contribution to the interval width ($2c$, Step 5 in Algorithm˜1) increases (Fig.˜2, middle and left, $\square$). The full confidence interval width, $2B+2c\Delta(\alpha)$, is not monotonic in $L$.
Tree Cover Linear Regression. We use a linear regression model $Y_{\text{Tree Cover \%}}=\sum_{p\in\mathcal{P}}\theta_{p}X_{p}$ to quantify how tree cover percentage in the contiguous United States (CONUS) in the year 2021 relates to three variables, $\mathcal{P}=\{\text{Aridity\_Index, Elevation, Slope}\}$. We use the 983 data points from Lu et al., [37], who in turn draw on [68, 65, 43]. We define our target region in the West portion of CONUS as locations with latitude in the range (25, 50) and longitude in the range (-125, -110). Out of all points in this region, we designate $50\%$ — totaling 133 sites — as target data. Next, we select the source data by taking a uniform random sample of $20\%$ of the remaining spatial locations, repeated over 250 random seeds to assess coverage performance. Each seed yields 170 source locations. Fig.˜7 illustrates the spatial split between source and target data for a representative seed. We discuss the data and our pre-processing in detail in Appendix˜E.
@@
We compare coverage for confidence intervals of the three parameters, $\theta_{\text{Aridity\_Index}},\theta_{\text{Elevation}},\text{ and }\theta_{\text{Slope}}$. We discuss how we evaluate coverage in Section˜E.3. In the top row of Fig.˜3, we see that our method is the only one to achieve the $95\%$ nominal coverage for all three parameters. Conversely, for the Slope parameter, every other method achieves coverage at most $54\%$. In the bottom row of Fig.˜10, we again see that alternative methods fail to provide coverage due to their overconfidence (small widths); see also Fig.˜10, which shows all methods’ constructed confidence intervals across all three parameters for 7 of the 250 seeds. In Section˜B.4, we further discuss how our intervals are the narrowest intervals among those that maintain validity. In Section˜E.5, we conduct a similar analysis but with target locations in the Southeast, rather than West, of CONUS. The results align with our discussion here.
@@
Choice of Lipschitz Constant in the Tree Cover Experiment. For the tree cover experiment, we leverage domain knowledge to set the Lipschitz constant to $L=0.2$, in units of percent tree cover per kilometer (km). This choice implies that a $1\%$ change in tree cover corresponds to moving $1/0.2=5$ km. To arrive at this choice, we observe that in certain regions of the U.S., such as the Midwest, tree coverage remains relatively uniform over several kilometers, so smaller Lipschitz constants (e.g., $L=0.02$, corresponding to a $1\%$ change over 50 km) would be appropriate. However, in other regions — such as the western U.S., where elevation changes are more pronounced (e.g., the Rockies, California, and the Pacific Northwest) — tree cover can change sharply over short distances. To account for these variations conservatively, we choose $L=0.2$. More generally, for real-world applications, we recommend the following strategy: (i) use domain knowledge to select a reasonable Lipschitz constant for the response variable, and (ii) inflate the Lipschitz constant to ensure a conservative estimate (which is more likely to satisfy Assumption˜4).
@@
Effect of Lipschitz Constant on Confidence Intervals in Tree Cover Experiment. In Fig.˜9, we show the coverage and width of our confidence intervals across three orders of magnitude of Lipschitz constants ($L$ from 0.001 to 1). For $L$ varying between 0.1 and 1 (a single order of magnitude variation around our chosen value of 0.2), we find that coverage is always met except in one case (Aridity Index with $L=0.1$), where it is very close to nominal (89% instead of 95%). For Slope and Elevation, coverage is met or very nearly met for all $L$ values. For Aridity Index, coverage is low for $L\leq 0.1$. Meanwhile, confidence intervals become noticeably wider for $L>0.5$ while remaining relatively stable for smaller values. These results support our intuition to err on the side of larger $L$ values to be conservative and maintain coverage.